OpenStudy (anonymous):

what is 2% of 145 and can you please show me how you got the answer-thanks

myininaya (myininaya):

.02*145

OpenStudy (anonymous):

2/100 = x/45

OpenStudy (anonymous):

2*(0.01*145)=2.9

myininaya (myininaya):

.02*145=2.9